# AerialBooth: Mutual Information Guidance for Text Controlled Aerial View Synthesis from a Single Image

### v1

@@ Abstract
We present a novel method, AerialBooth, for synthesizing the aerial view from a single input image using its text description. We leverage the pretrained text-to-2D image stable diffusion model as prior knowledge of the 3D world. The model is finetuned in two steps to optimize for the text embedding and the UNet that reconstruct the input image and its inverse perspective mapping respectively. The inverse perspective mapping creates variance within the text-image space of the diffusion model, while providing weak guidance for aerial view synthesis. At inference, we steer the contents of the generated image towards the input image using novel mutual information guidance that maximizes the information content between the probability distributions of the two images. We evaluate our approach on a wide spectrum of real and synthetic data, including natural scenes, indoor scenes, human action, etc. Through extensive experiments and ablation studies, we demonstrate the effectiveness of AerialBooth and also its generalizability to other text-controlled views. We also show that AerialBooth achieves the best viewpoint-fidelity trade-off though quantitative evaluation on $7$ metrics analyzing viewpoint and fidelity w.r.t. input image. Code and data is available at https://github.com/divyakraman/AerialBooth2023.
Figure 1: Given a single input image and its text description, AerialBooth generates an aerial view of the given image. It leverages pretrained text to 2D image models and does not require access to any paired data or explicit 3D or multi-view information during training or test time.
@@ 1 Introduction
Widely available 2D text-to-image models such as Stable Diffusion [49], trained on large-scale text-image data, contain immense knowledge of the 3D world. They are capable of generating various viewpoints of objects. Consequently, it can be inferred that text, along with 2D images, form a rich representation of the 3D space. However, given an in-the-wild image and its text description, the vanilla stable diffusion model may not be able to generate different viewpoints of the scene in the image due to limitations of the expressiveness of the text. Image personalization and editing approaches such as DreamBooth [50], Imagic [25] are successful in performing a wide range of non-rigid transformations including text-controlled view synthesis. However, the large translation required for novel view synthesis, especially aerial view, makes them sub-optimal [42] in generating the aerial viewpoint due to bias-variance tradeoff issues, especially when a single input image is provided. In this paper, we focus on text controlled aerial view synthesis from a single input image.
The usage of text to control the viewpoint makes the definition of ‘aerial view’ ambiguous since no camera parameters are explicitly involved. Without loss of generality, we consider any generated image that has a higher camera angle than the input image to be an aerial view of the image. The task of generating aerial views from a given image, also termed as cross-view synthesis, is an important computer vision task with applications [36] in cross-view localization and mapping [22], autonomous driving [7], augmented and virtual reality [15], 3D reconstruction [67], medical imaging [65], security and surveillance [2] and generation of realistic aerial view synthetic data for improved aerial view perception tasks. However, it is difficult to collect paired data and 3D metadata to train large generative models [42]. Consequentially, it is beneficial to develop an unsupervised method, that doesn’t use multi-view or depth information at train or test time, to generate aerial views from a single input image.
Prior work on text-controlled aerial view synthesis such as Aerial Diffusion [27] doesn’t work too well on complex scenes such as traffic scenes, indoor scenes and scenes with multiple objects and requires hyperparameter tuning for every input image. Moreover, many of the resulting images have residual diagonal artifacts. A closely related problem is novel view synthesis (NVS) and 3D synthesis [35, 42]. NVS methods sample different camera viewpoints by explicitly specifying the camera angle. However, they often need to trained on an enormous amount of large-scale datasets with scenes from different views and 3D data.
@@
We propose AerialBooth for aerial view synthesis from a single image. Our method leverages text-to-image diffusion models for prior knowledge and does not require any 3D or multi-view data. Moreover, our method performs test-time finetuning and hence does not require any additional training, it is easily generalizable to any in-the-wild image. The key novel components of our algorithm include:
@@
    Test-time finetuning: We first prepare the text-image embedding space by sequentially optimizing the text embedding and the diffusion model on the input image and its inverse perspective mapping in its close vicinity. Such a mechanism enables the model to learn the characteristics of the input image while acquiring adequate variance to generate aerial views.
@@
    Mutual Information Guided Inference: To generate the aerial view of the image using its text description, we propose a mutual information guidance formulation. Mutual information guidance steers the content in the generated aerial view towards the input image by maximizing the information content between the probability distributions of the input image and the generated aerial image, without imposing any pixel-to-pixel matching constraints between the two images with vastly different viewpoints.
Evaluation: To test our method, we collect a diverse set of synthetic images (from Stable Diffusion XL) and real images (from Unsplash), spanning from natural scenes, indoor scenes, to human actions and animations. Comparisons with prior work on single-image text-based view synthesis show that AerialBooth achieves a better viewpoint-fidelity (w.r.t. input image) trade-off to generate high-quality aerial images of the given input images, as demonstrated by our qualitative as well as quantitative results on $7$ metrics including CLIP Score [44], SSCD Score [40] and CLIP Directional Similarity [14]. We conduct extensive ablation experiments to show the usefulness of the inverse perpective mapping is useful in generating aerial views of images, and benefits of mutual information in guiding the contents in the generated image towards the contents of the input image. Our method can also be extended to generate more views that can be text-controlled (such as ‘side view’, bottom view’, ‘back view’), as evidenced by our results.
@@ 2 Related work
Novel view synthesis [68, 64, 39] from a single image is an active area of research in generative AI. Many methods [60, 24, 17, 75] use NeRF based techniques. Nerdi [11] use language guidance with NeRFs for view synthesis. Many recent methods use diffusion [34, 57, 35, 43, 58, 55, 5, 52] to sample different views. 3D generation methods [42, 30, 71, 45, 8] use text to guide the reconstruction. All of these methods use large amounts of multi-view and 3D data for supervised training. Methods like Zero-1-to-3 [35] and Zero-123++ [55] use a pretrained stable diffusion [49] model, along with large data for supervised training, to learn different camera viewpoints. In contrast, we use the pretrained stable diffusion model, and don’t use any multi-view or 3D or any other kind of training data. Our method performs text-controlled view synthesis on a single input image, this paper doesn’t focus on explicitly controlling the (camera) angle of the generated view, the generated image is of arbitrary view, within the limits of the text guidance for viewpoint.
Image editing/ personalization.
@@
Diffusion models have emerged as successful tools for single image editing and personalization. Methods such as DreamBooth [50], DreamBooth LoRA [21], HyperDreamBooth [51], Textual Inversion [16], Custom Diffusion [28] are able to generated personalized images of subjects. Image editing methods such as Imagic [25], Paint-by-Example [72], ControlNet [73], DiffEdit [9] are able to edit images to perform non-rigid transformations and also use exemplar signals for guidance. However, these methods can either generate aerial images with low fidelity w.r.t. the input image or generate high-fidelity images with viewpoints very close to the input image. In contrast, our method is able to execute larger view transformations than prior work, while preserving fidelity w.r.t. the input image.
Cross-view synthesis.
Prior work on cross-view synthesis [46, 61, 48, 63, 13, 36, 32, 56, 31, 47, 33, 69, 54, 1, 74] use paired data and modalities such as semantic maps, depth, multi-views, etc within their architectures. In this paper, we propose an unsupervised method for view synthesis using pretrained 2D text to image diffusion models. Aerial Diffusion [27] trains on homography images corresponding to Inverse Perspective Mapping (IPM) to generate aerial images by alternatingly sampling between prompts corresponding to viewpoint and the homography image. However, the diagonal artifacts in the resulting images indicate a need to utilize IPM better. Moreover, the quality of the generated images is very poor for complex scenes that typically contain more than one object. Our method is capable of generating high quality images for diverse complex scenes.
Guidance techniques in diffusion.
Guidance methods [19, 3, 12, 38] have been used to control and guide diffusion denoising towards semantic maps, image classes, etc. In this paper, we propose a novel diffusion guidance formulation to align the contents in two images that have vastly different views.
@@
Figure 2: Overview. AerialBooth takes in a single input image $I_{S}$ to generate its aerial view using a pretrained text-to-image diffusion model, and without any 3D or multi-view supervision. It finetunes the diffusion model in two stages to optimize the text embedding and the diffusion model to reconstruct the input image and its inverse perspective mapping (or homography image) in close vicinity. Such a mechanism enables the incorporation of image specific knowledge within the model, while retaining its imaginative capabilities (or variance). At inference, AerialBooth uses mutual information guidance to maximize the information between the probability distributions of the generated image and $I_{S}$, to generate a high-fidelity image.
@@ 3 Method
We present AerialBooth  to generate aerial view images from a single input image $I_{S}$ and its text description $t_{S}$ (e.g. ‘a cosy living room’). The text description of the image can be obtained using the BLIP-2 model [29]. We leverage the pretrained text-to-2D image stable diffusion [49] model to serve as a strong prior. Furthermore, we do not use any training data or 3D/multi-view data. We present an overview of our method in Figure 2.
  • Finetuning the diffusion model: We perform a two-stage test-time finetuning on the input image $I_{S}$ to incorporate the knowledge pertaining to the input image within the pretrained model, while ensuring that the model does not get biased towards it. Specifically, we optimize the text embedding as well as the diffusion UNet sequentially on the input image and its inverse perspective mapping, in close vicinity.
  • Inference: To generate the aerial-view image, we use the target text description $t_{T}$ that takes the form ‘aerial view, ’ + $t_{S}$ (e.g. ‘aerial view, a cosy living room’). To ensure that the generated image has high fidelity with the input image, we use mutual information guidance.
@@
3.1 Test-time finetuning
The text-to-2D image stable diffusion model has knowledge of the 3D world as a consequence of the large amount of diverse data it has been trained on. It understands [53] different viewpoints, different styles, backgrounds, etc. Image editing and personalization methods such as DreamBooth [50], DreamBooth LoRA [21], Imagic [25], SVDiff [18] exploit this property to perform transformations such as making a standing dog sit and generating its image in front of the Eiffel tower. At a high level, the standard procedure adopted by these methods to generate edited or personalized images is to finetune the model, followed by inferencing. These methods are however not very successful in text-guided aerial view synthesis, which demands a large transformation. Specifically, directly finetuning the diffusion model on $e_{S}$ to reconstruct $I_{S}$ results in severe overfitting, where $e_{S}$ is the CLIP text embedding for $t_{S}$. This makes it difficult for the model to generate large variations to the scene required for aerial view synthesis.
We propose a two-step finetuning approach to enable the model to learn the characteristics of $I_{S}$, while ensuring sufficient variance for aerial view generation.
3.1.1 Optimized Text Embedding:
In the first step, we start from $e_{S}$ and compute the optimized text embedding $e_{opt}$ to reconstruct $I_{S}$ using a frozen diffusion model UNet using the denoising diffusion loss function $L$ [20].
@@ 3.1.1 Optimized Text Embedding:
where $t$ is the diffusion timestep and $x_{t}$ is the latents at time $t$. This formulation allows us to find the text embedding that characterizes $I_{S}$ better than the generic text embedding $e_{S}$. Next, to enable $e_{opt}$ accurately reconstruct $I_{S}$, we optimize the diffusion UNet using the denoising diffusion objective function. Note that we insert LoRA layers within the attention modules in the diffusion UNet and finetune only the LoRA layers with parameters $\theta_{LoRA}$, the rest of the UNet parameters are frozen,
@@
3.1.2 Inverse Perspective Mapping:
Inverse perspective mapping (IPM) [59] is a homography transformation from classical computer vision to generate the aerial-view of an image from its ground-view. Despite not being accurate, it can provide pseudo weak supervision for the generation of the aerial image and also add more variance to the embedding space. We denote the inverse perspective mapping of the input image by $I_{H}$, computed following [27]. In the second step of finetuning, to find the text embedding $e_{H}$ that best characterises $I_{H}$, we start from $e_{opt}$ and optimize the text embedding with a frozen diffusion model, similar to Equation 1. Finding $e_{H}$ in the vicinity of $e_{opt}$ instead of $e_{S}$ ensures that the text-image space corresponding to $e_{S}$ doesn’t get distorted to generate the poor quality IPM image. Next, we finetune the diffusion model using the denoising diffusion objective function to reconstruct $I_{H}$ at $e_{H}$, similar to Equation 2. Again, only the LoRA layers are finetuned, the rest of the UNet is frozen.
Note that we find $e_{opt}$ and $e_{H}$ by optimizing $e_{S}$ and $e_{opt}$, respectively for a small number of iterations. We need to ensure that $e_{S}$, $e_{opt}$ and $e_{H}$ are all in close vicinity. Our finetuning approach conditions the embedding space to encapsulate the details of $I_{S}$, while being able to generate large transformations required for the generation of the aerial image.
3.2 Inference
Our next step is to use the finetuned diffusion model to generate the aerial view image for the text prompt $t_{T}$. The text embedding for $t_{T}$ is $e_{T}$. Diffusion denoising, conditioned on $e_{T}$ is capable of generating aerial images corresponding to $I_{S}$. However, oftentimes, the contents of the generated aerial view image does not completely align with the contents of $I_{S}$. Consequently, to ensure high fidelity generations, our goal is to guide the contents of the aerial view image towards the contents of $I_{S}$.

### v3

# HawkI: Homography & Mutual Information Guidance for 3D-free Single Image to Aerial View

@@ Abstract
We present HawkI, for synthesizing aerial-view images from text and an exemplar image, without any additional multi-view or 3D information for finetuning or at inference. HawkI uses techniques from classical computer vision and information theory. It seamlessly blends the visual features from the input image within a pretrained text-to-2D-image stable diffusion model with a test-time optimization process for a careful bias-variance trade-off, which uses an Inverse Perspective Mapping (IPM) homography transformation to provide subtle cues for aerial-view synthesis. At inference, HawkI employs a unique mutual information guidance formulation to steer the generated image towards faithfully replicating the semantic details of the input-image, while maintaining a realistic aerial perspective. Mutual information guidance maximizes the semantic consistency between the generated image and the input image, without enforcing pixel-level correspondence between vastly different viewpoints. Through extensive qualitative and quantitative comparisons against text + exemplar-image based methods and 3D/ multi-view based novel-view synthesis methods on proposed synthetic and real datasets, we demonstrate that our method achieves a significantly better bias-variance trade-off towards generating high fidelity aerial-view images. Code and data is available at https://github.com/divyakraman/HawkI2024.
Figure 1: HawkI generates aerial-view images from a text description and an exemplar input image. It builds on a text to 2D image stable diffusion model and does not require any additional 3D or multi-view information at fine-tuning or inference.
@@ 1 Introduction
Widely available text-to-image models such as Stable Diffusion [64], trained on large-scale text and 2D image data, contain rich knowledge of the 3D world. They are capable of generating scenes from various viewpoints, including aerial views. However, due to limitations of the expressiveness of the text, we may not be able to completely describe the precise scene that we wish to generate. Moreover, the generative capabilities of the pretrained model is constrained by the aerial-view images in the dataset that it was trained on, which is typically limited. Consequentially, along with text, it is beneficial to use an easily available representative front-view image describing the aerial view of the scene we wish to generate. The task of generating aerial-view images from a given input image and its text description finds applications in the generation of realistic diverse aerial view synthetic data for improved aerial view perception tasks [36, 41, 37, 12, 4], and weak supervision for cross-view synthesis applications [49] such as localization and mapping [31], autonomous driving [10], augmented and virtual reality [21], 3D reconstruction [82], medical imaging [80], drone-enabled surveillance [2].
Aerial-view images corresponding to text and an input image can be sampled using text-to-3D and novel view synthesis (NVS) [48, 56]. These methods sample different camera viewpoints by explicitly specifying the camera angle. However, they often need to trained on enormous, large-scale datasets with 3D details and scenes from multiple views. Is it possible for text-to-image(2D) diffusion models to generate aerial-view images without any multi-view or 3D information?
Another closely related task is image editing [34] and personalization [65], where the goal is to use an input image and a target text to generate an image consistent with both inputs. These methods are generally successful in performing a wide range of non-rigid transformations including text-controlled view synthesis. However, the large translation required for aerial view synthesis makes them sub-optimal [56]. This is due to bias-variance trade-off issues, even more amplified when only a single input image is provided. Aerial Diffusion [38] attempted to alleviate this bias-variance trade-off, but at the cost of per-sample hyperparameter tuning, residual diagonal artifacts in many of the generated images arising from direct finetuning on sub-optimal homography projections and severe performance drops on complex scenes with multiple objects.
@@
We propose HawkI for aerial view synthesis, guided by text and a single input image. Our method leverages text-to-image diffusion models for prior knowledge and does not require any 3D or multi-view data. Since explicitly specifying camera details in text descriptions isn’t always possible, similar to prior work on text-based viewpoint generation [65, 38], we consider any generated image with a significantly higher viewpoint and altitude compared to the original image to be an aerial view. HawkI fuses techniques from classical computer vision and information theory within a stable diffusion backbone model to guide the synthesis of the aerial-view image. The key novel components of our algorithm include:
@@
    Test-time optimization: This step enables the model to acquire the characteristics of the input image, while maintaining sufficient variability in the embedding space for aerial-view synthesis. We condition the embedding space by sequentially optimizing the CLIP text-image embedding and the LoRA layers corresponding to the diffusion UNet on the input image and its Inverse Perspective Mapping (IPM) homography transformation in close vicinity. In addition to creating variance, IPM provides implicit guidance towards the direction of transformation for aerial-view synthesis.
@@
    Mutual Information Guided Inference: This step generates a semantically consistent aerial-view image while accounting for viewpoint differences. Unlike conventional approaches [3, 22] that rely on restrictive pixel-level constraints (often ineffective for different viewpoints), we propose a mutual information vastly guidance formulation. Mutual information guidance, rooted in information theory, ensures consistency between the contents of the generated image and the input image by maximizing the information contained between the probability distributions of the input image and the generated aerial image.
Our method performs inference-time optimization on the given text-image inputs and does not require a dataset to train on, hence, it is easily applicable to any in-the-wild image. To test our method, we collect a diverse set of synthetic images (from Stable Diffusion XL) and real images (from Unsplash), spanning across natural scenes, indoor scenes, human actions and animations. Qualitative and quantitative comparisons with prior work, on metrics such as CLIP [58] (measuring viewpoint and text consistency) and SSCD [54], DINOv2 [52] (measuring consistency w.r.t. input image), demonstrate that HawkI generates aerial-view images with a significantly better viewpoint-fidelity (or bias-variance) trade-off. We also present extensive ablation experiments and comparisons with 3D-based novel-view synthesis methods highlighting the benefits of our 3D-free classical guidance approaches. Our method can also be extended to generate more views that can be text-controlled (such as ‘side view’, bottom view’, ‘back view’), as evidenced by our results.
@@ 2 Related work
3D and novel view synthesis. Novel view synthesis [83, 79, 53] from a single image is an active area of research in generative AI. Many methods [75, 33, 25, 92] use NeRF based techniques. Nerdi [16] use language guidance with NeRFs for view synthesis. Many recent methods use diffusion [47, 72, 48, 57, 73, 70, 6, 67] to sample different views. 3D generation methods [56, 42, 86, 59, 11] use text to guide the reconstruction. All of these methods use large amounts of multi-view and 3D data for supervised training. Methods like Zero-1-to-3 [48] and Zero-123++ [70] use a pretrained stable diffusion [64] model, along with large data for supervised training, to learn different camera viewpoints. 3D-free methods such as Free3D [91] still require multi-view and 3D information while training.
Warping, scene extrapolation and homography. Scenescape [23], DiffDreamer [7] and similar methods [83, 63, 9] estimate a depth map, reproject the pixels into the desired camera perspective and outpaint the scene. Again, these methods require 3D and multi-view information at training stage. Using a homography to estimate the scene from an aerial perspective is highly inaccurate, hence, attempting to create realistic aerial view images by simply filling in missing information based on the homography (outpainting) leads to poor outcomes. Homography maps have also been used in various deep learning based computer vision solutions [88, 18, 46, 26, 14].
@@
Image editing/ personalization. Diffusion models have emerged as successful tools for single image editing and personalization. Methods such as DreamBooth [65], DreamBooth LoRA [30], HyperDreamBooth [66], Textual Inversion [24], Custom Diffusion [39] are able to generated personalized images of subjects. Image editing methods such as Imagic [34], Paint-by-Example [87], ControlNet [89], DiffEdit [13] are able to edit images to perform non-rigid transformations and also use exemplar signals for guidance. However, these methods can either generate aerial images with low fidelity w.r.t. the input image or generate high-fidelity images with viewpoints very close to the input image.
@@
Figure 2: Overview. HawkI generates aerial-view images, using a text description and a single image $I_{S}$ as supervisory signals. It builds on a pretrained text-to-image diffusion model, and does not use any 3D or multi-view information. It performs test-time finetuning to optimize the text embedding and the diffusion model to reconstruct the input image and its inverse perspective mapping in close vicinity. Such a mechanism enables the incorporation of image specific knowledge within the model, while retaining its imaginative capabilities (or variance). At inference, HawkI uses mutual information guidance to maximize the information between the probability distributions of the generated image and $I_{S}$, to generate a high-fidelity aerial-view image.
Cross-view synthesis. Prior work on cross-view synthesis [60, 76, 62, 78, 19, 49, 45, 71, 44, 61, 43, 84, 69, 1, 90] are data intensive - they use paired data and modalities such as semantic maps, depth, multi-views, etc within their architectures. Aerial Diffusion [38] uses text and an exemplar image for the task by alternating sampling between viewpoint and homography projects. However, the generated images have diagonal artifacts with poor quality results for complex scenes that typically contain more than one object and requires manual per-sample hyperparameter tuning.
Guidance techniques in diffusion. Guidance methods [29, 3, 17, 51] have been used to control and guide diffusion denoising towards semantic maps, image classes, etc. These guidance techniques cannot enforce view-invariant image-image similarity, critical for aligning the contents in two images with vastly different viewpoints.
@@ 3 Method
We present HawkI  to generate aerial view images using a single input image $I_{S}$ and its text description $t_{S}$ (e.g. ‘a cosy living room’, can be obtained using the BLIP-2 model [40]). We do not use any training data or 3D/multi-view data. We leverage the pretrained text-to-2D image stable diffusion [64] model to serve as a strong prior, and utilize classical computer vision and information theory principles to achieve the desired goal in a holistic manner. We present an overview of our method in Figure 2.
  • Test-time optimization: We perform multi-step test-time optimization to incorporate the input image $I_{S}$ within the pretrained model, at an appropriate bias-variance trade-off. Specifically, we optimize the CLIP text-image embedding and the LoRA layers in the diffusion UNet sequentially on the input image and its inverse perspective mapping, in close vicinity. This additionally conditions the embedding space viewpoint transformations, along with acquiring the characteristics of the input image.
  • Inference: To generate the aerial-view image, we use the target text description $t_{T}$, that takes the form ‘aerial view, ’ + $t_{S}$ (e.g. ‘aerial view, a cosy living room’). To ensure that the generated aerial image is semantically close to the input image, we use mutual information guidance.
@@
3.1 Test-time optimization
The text-to-2D image stable diffusion model has knowledge of the 3D world as a consequence of the large amount of diverse data it has been trained on. It understands [68] different viewpoints, different styles, backgrounds, etc. Image editing and personalization methods such as DreamBooth [65], DreamBooth LoRA [30], Imagic [34], SVDiff [27] exploit this property to perform transformations such as making a standing dog sit and generating its image in front of the Eiffel tower. At a high level, the standard procedure adopted by these methods to generate edited or personalized images is to finetune the model on the input image, followed by inferencing. These methods are however not very successful in text-guided aerial view synthesis, which demands a large transformation. Specifically, directly finetuning the diffusion model on $e_{S}$ to reconstruct $I_{S}$ results in severe overfitting, where $e_{S}$ is the CLIP text embedding for $t_{S}$. This makes it difficult for the model to generate large variations to the scene required for aerial view synthesis.
We propose a four-step finetuning approach to enable the model to learn the characteristics of $I_{S}$, while ensuring sufficient variance for aerial view generation.
3.1.1 Optimization using $I_{S}$:
In the first step, we start from $e_{S}$ and compute the optimized CLIP text-image embedding $e_{opt}$ to reconstruct $I_{S}$ using a frozen diffusion model UNet using the denoising diffusion loss function $L$ [28].
@@ 3.1.1 Optimization using $I_{S}$:
where $t$ is the diffusion timestep and $x_{t}$ is the latents at time $t$. This formulation allows us to find the text embedding that characterizes $I_{S}$ better than the generic text embedding $e_{S}$.
Next, to enable $e_{opt}$ accurately reconstruct $I_{S}$, we optimize the diffusion UNet using the denoising diffusion objective function. Note that we insert LoRA layers within the attention modules in the diffusion UNet and finetune only the LoRA layers with parameters $\theta_{LoRA}$, the rest of the UNet parameters are frozen,
@@
3.1.2 Optimization using inverse perspective mapping.
Inverse perspective mapping (IPM) [74] is a homography transformation from classical computer vision to generate the aerial-view of an image from its ground-view. Despite not being accurate, it can provide pseudo weak supervision for the generation of the aerial image and also add more variance to the embedding space. We denote the inverse perspective mapping of the input image by $I_{H}$, computed following [38]. We perform the following optimization steps to condition the embedding space towards the desired viewpoint transformation. To find the text embedding $e_{H}$ that best characterises $I_{H}$, we start from $e_{opt}$ and optimize the text embedding with a frozen diffusion model, similar to Equation 1. Finding $e_{H}$ in the vicinity of $e_{opt}$ instead of $e_{S}$ ensures that the text-image space corresponding to $e_{S}$ doesn’t get distorted to generate the poor quality IPM image. Next, we finetune the diffusion model using the denoising diffusion objective function to reconstruct $I_{H}$ at $e_{H}$, similar to Equation 2. Again, only the LoRA layers are finetuned, the rest of the UNet is frozen.
Note that we find $e_{opt}$ and $e_{H}$ by optimizing $e_{S}$ and $e_{opt}$, respectively for a small number of iterations. We need to ensure that $e_{S}$, $e_{opt}$ and $e_{H}$ are all in close vicinity. Our finetuning approach conditions the embedding space to encapsulate the details of $I_{S}$ and viewpoint, while having sufficient variance to generate large transformations required for the generation of the aerial image.
3.2 Mutual Information Guided Inference
Our next step is to use the finetuned diffusion model to generate the aerial view image for the text prompt $t_{T}$. The text embedding for $t_{T}$ is $e_{T}$. Diffusion denoising, conditioned on $e_{T}$ is capable of generating aerial images corresponding to $I_{S}$. However, oftentimes, the contents of the generated aerial view image does not align well with the contents of $I_{S}$. Consequently, to ensure high fidelity generations, our goal is to guide the contents of the aerial view image towards the contents of $I_{S}$.